The number of moles of a sodium salt (A) is defined as m = ∫01 [1/1+x2 ]dx. Given that the sodium salt (A) reacts with concentrated sulphuric acid in the ratio 2:1 to produce a product B according to the equation


A + H2SO4 → B. Find the mass of the compound B that was produced (RMM of B= 20g/mol)


A. (π/4)g B. (20.5 +π/4)g


C. (5/3)m2 D. (5/2)π


1
Expert's answer
2021-08-10T05:03:31-0400

"m = ∫^1_0 \dfrac{1}{1+x² } \ dx"


"m = ∫^1_0 \dfrac{1}{1+x² } \ dx =\dfracπ4"


"2NaCl (A)+ H_2SO_4 \to Na_2SO_4 + 2HCl\qquad(\approx B)"

ratio of moles = "2:1"

since the number of moles of A is "\dfracπ4", then the number of moles of B is "\dfracπ4÷2 =\dfracπ8"


mass = number of moles × mass = "\dfracπ8×20 = \dfrac{5\pi}{2}"

Ans. d.